OpenStudy (anonymous):

A bag contains 45 marbles of various colors. Lena draws 48 times, each time replacing the marble before drawing again. She gets a total of 16 red marbles. Which is best prediction of how many red marbles are in the bag? A. 3 B. 13 C. 15 D. 16 plz help i am on a test

OpenStudy (anonymous):

can anyone plz help

OpenStudy (dumbcow):

use a proportion \[\frac{R}{45} = \frac{16}{48}\]

OpenStudy (anonymous):

thats the thing i dont get it

OpenStudy (dumbcow):

btw dont say you are on a test....that would imply you are cheating:{

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

solve proportions by reducing the fraction 16/48 = 1/3

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

get "r" by itself by multiplying both sides by 45

OpenStudy (anonymous):

so 16 time 45 and 49 time 45

OpenStudy (dumbcow):

hmm... 16/48 = 1/3 .... 16*3 = 48 right?

OpenStudy (dumbcow):

\[\frac{R}{45} = \frac{1}{3}\] \[\frac{45}{45} R = \frac{45}{3}\] R = 45/3